An ideal vapor compression refrigeration cycle uses refrigerant 12 as a working fluid in an air conditioning system. The refrigerant enter the compressor as saturated vapor at 5oC and leaves the condenser as saturated liquid at 55oC. The mass flow rate of refrigerant is 0.7 kg/s. Heat is transferred from a reservoir at 15oC (the cool space) to the refrigerant in the evaporator and the heat rejected by the condenser is transferred to the environment which is at a temperature of 45oC

  1. Calculate the irreversibility for each component and for the cycle
  2. Calculate the relative efficiency or utilization factor for the refrigeration cycle which is defined as the ratio of minimum exergy requirement divided by actual exergy input.
Answer #1

for compressor 19 cr for Condenfor TR2 = mo (S₂-1) = 0 PRys MiTo [CS-s) + Pol M = (ho-ho) =(45+271) [(0.9 191.42 -90.29 3196